An oil tank is 5/8 full. When 6 gallons of oil are removed, the tank is 1/4 full. What is the total capacity of the tank in gallons?
Found 2 solutions by Cromlix, stanbon:
Answer by Cromlix(4381)

5/8 - 1/4 = 3/8
3/8 = 6 gallons
8/8 = 16 gallons
Hope this helps.
:-)

Answer by stanbon(75887)

An oil tank is 5/8 full. When 6 gallons of oil are removed, the tank is 1/4 full. What is the total capacity of the tank in gallons?
-------
Let total capacity be "x"::
Equation:
(5/8)x - 6 = (1/4)x
----
Multiply thru by 8 to get:
5x - 48 = 2x
3x = 48
x = 16 gallons
----------------------
Cheers,
Stan H.
--------------